1. Write the equation of the curve for which y'''=-96x, if the curve is tangent to the line 20x+y+1=0 at (1, -3) and passes thru (0, 1)?


2. A particle moves such that its acceleration is defined by the equation a=3t²-2. When t=0, s=5 and when t=1, v=-20. What is s when t=2?


3. A sandbag is dropped from a balloon rising at 9.6 m/s at a height of 64 m. Find the highest point reached by the sandbag and the time of flight. Note, use g=-10 m/s².


1
Expert's answer
2021-05-12T06:53:45-0400

Ans:-"\Rightarrow a=3t²-2"

"\dfrac{dv}{dt}=3t^2-2\\"

integrating with respect to t

"v=t^3-2t+c_1"

at

"t=1, v=-20 \to c_1=-19\\
\Rightarrow v=t^3-2t-19"

"\Rightarrow \dfrac{ds}{dt}=t^3-2t-19"

again integrate with respect to t

"\Rightarrow s=\dfrac{t^4}{4}-t-19t+c_2"

at "t=0, s=5 \to c_2 =5"

then required equation for "s=\dfrac{t^4}{4}-t-19t+5\\"

then at "t=2 ,s=-31"

(3) Ans:-

"\Rightarrow x=ut+\dfrac{1}{2}at^2\\
\Rightarrow -64=9.6+\dfrac{1}{2}\times (-10)\times t^2\\
\Rightarrow 5t^2-9.6-64=0\\
\Rightarrow t=4.66sec"

So, the time of flight of sandbag will be "4.66 sec"

Displacement of balloon after "4.66 sec =9.6*4.66=44.736" m

Balloon was at 64m height when sandbag was dropped

Highest height delivered by the sandbag will be"64+44.736=108.736m"

(1)Ans:-

"y'''=-96x\\"

integrate with respect to x

"\dfrac{d^2y}{dx^2}=-48x^2+c_1\\"


"\dfrac{dy}{dx}=-16x^3+c_1x+c_2\\"


we know that the curve is tangent to the line "y=-20x-1" then

"\dfrac{dy}{dx}=-20=-16x^3+c_1x+c_2"

then put"(0,1)" and "(1,-3)" then equation will be

"\dfrac{dy}{dx}=-16x^3+16x-20"

again integrate with respect to "x"

"y=4x^4+8x^2-20x"